OpenStudy (anonymous):

Vivian combined different amounts of white paint, blue paint, and green paint to make 145 milliliters (ml) of paint of a new shade. The amount of white paint was x ml. The amount of blue paint was 10 ml less than twice the amount of white paint, and the amount of green paint was half the amount of blue paint. This is represented in the equation below. x + (2x – 10) + (2x – 10) = 145

OpenStudy (anonymous):

Just a couple things, futrchamps solution does not add up to the 145 that the equation calls for, if you plug in 25 for x in the equation it adds to 25+(50-10)+(50-10) or 25+40+40 which equals 105 not 145. Secondly, the equation in the problem doesn't match the story problem since it says that the green paint is half the amount of the blue paint. So in my solution, I will give two answers, one to the equation in the problem, and one to the word problem itself. First, the equation: x+(2x-10)+(2x-10)=145 simplified to: 5x-20=145 5x=165 x=33 plugging that back into the equation yields: 33+(66-10)+(66-10)=145 However, according to this equation, blue and green paints would have the same amount while the story problem said that green was half that amount. As for the story problem, it would actually look something like this: x+(2x-10)+((2x-10)/2)=145 This problem is a little more confusing since it brings division and an extra set of parentheses. You can still simplify it though: 3x-10+(x-5)=145 4x-15=145 4x=160 x=40 Plug it back in: 40+(80-10)+((80-10)/2) 40+70+35=145 Keep in mind that x in this equation is the amount of white paint, the first parentheses is the blue paint, and the third set is the green.